# Generalized Frobenius Manifold Structures on the Orbit Spaces of Affine Weyl Groups I

### v1

@@ Abstract.
We present an approach to construct a class of generalized Frobenius manifold structures on the orbit spaces of affine Weyl groups, and prove that their monodromy groups are proper subgroups of the associated affine Weyl groups.
@@
Generalized Frobenius manifold, Affine Weyl group, Flat pencil of metrics, Monodromy Group
@@ 1. Introduction
The notion of Frobenius manifold was introduced by Dubrovin in the early 1990s in his study of the relation of integrable systems and 2-dimensional topological field theories [8, 9]. This geometric structure and its generalizations play a central role in topological field theory, Gromov-Witten theory, singularity theory and in the theory of integrable systems (see [1, 2, 3, 4, 5, 9, 13, 15, 18, 19, 22, 24] and the references therein).
@@
We consider in this paper the construction of a class of generalized Frobenius manifold structures on the orbit spaces of affine Weyl groups. These structures satisfy all the axioms of Dubrovin’s definition of Frobenius manifolds except for the flatness of the unit vector field. The simplest example in this class corresponds to the affine Weyl group of type $A_{1}$, which naturally appears in the study of Hodge integrals and their remarkable connections to two important discrete integrable systems: the Volterra hierarchy and the $q$-deformed KdV hierarchy—both fundamental in soliton theory [21, 22].
@@
Our approach parallels the construction of a Frobenius manifold structure on the orbit space of a Coxeter group or on a certain extension of an affine Weyl group [9, 12, 14, 25, 26, 28, 29]. For these cases, a flat pencil of metrics, based on which the Frobenius manifold structure is constructed, is obtained by projecting the Euclidean metric (on the space where the group acts) onto the orbit space. The key to these constructions lies in the linear dependence of both the contravariant components of the metric and its Levi-Civita connection on a distinguished coordinate of the orbit space—this special coordinate naturally gives rise to the flat unit vector field of the Frobenius manifold. However, for the generalized Frobenius manifolds that we are considering, the unit vector field is not flat, implying the absence of such a special coordinate. To overcome this, we proceed as follows: Given an affine Weyl group $W_{a}$ acting on an $\ell$-dimensional Euclidean space $V$, we introduce a parameter $\lambda$ into the invariant Fourier polynomials, yielding the coordinates $z^{1},\dots,z^{\ell}$ on the orbit space. The projected contravariant metric $\left(g^{\alpha\beta}(z)\right)$ then depends on $\lambda$. The key idea is to choose the $\lambda$-dependence of the invariant Fourier polynomials such that $\left(g^{\alpha\beta}(z)\right)$ varies linearly with $\lambda$.
@@
In the present paper, we detail this construction, study properties of the resulting generalized Frobenius manifolds (including their monodromy groups), and provide some explicit examples. Further examples will appear in a subsequent paper [20]. The paper is organized as follows. In Sect. 2, we introduce the invariant $\lambda$-Fourier polynomial ring and study its properties. In Sect. 3, we construct generalized Frobenius manifold structures on orbit spaces of affine Weyl groups, and summarize the main results in Theorem 3.9. In Sect. 4, we analyze properties of the monodromy groups of these generalized Frobenius manifolds. In Sect. 5, we present some representative examples. In Sect. 6, we conclude the paper with remarks and further directions.
@@ 2.1. Affine coordinates
[…]
@@
where $c\in\mathbb{R}$ is a fixed parameter. We will also use the following expressions of elements of $V$:
@@
$\mathbf{x}\mapsto\sigma(\mathbf{x})+\sum_{j=1}^{\ell}k_{j}\alpha_{j}^{\vee},\quad\sigma\in W,\ k_{1},\dots,k_{\ell}\in\mathbb{Z},$ (2.5)
@@
The action
$\sigma\left(c\omega+x^{1}\alpha_{1}^{\vee}+\dots+x^{\ell}\alpha_{\ell}^{\vee}\right)=c\omega+\tilde{x}^{1}\alpha_{1}^{\vee}+\dots+\tilde{x}^{\ell}\alpha_{\ell}^{\vee}$ (2.6)
of $\sigma\in W_{a}(R)$ on $V$ yields the change of affine coordinates
$\sigma(x^{1},\dots,x^{\ell};c)=(\tilde{x}^{1},\dots,\tilde{x}^{\ell};c).$
2.2. The Invariant $\lambda$-Fourier polynomial ring and its generators
Let us consider the Fourier function ring
$\mathscr{F}=\mathrm{span}_{\mathbb{C}}\left\{m(x^{1},\dots,x^{\ell};c)=\mathrm{e}^{2\pi i(t_{0}c+t_{1}x^{1}+\dots+t_{\ell}x^{\ell})}\mid t_{0},t_{1},\dots,t_{\ell}\in\mathbb{R}\right\}.$
We introduce a gradation on $\mathscr{F}$ by the assignment
@@ 2.2. The Invariant $\lambda$-Fourier polynomial ring and its generators
where $\theta_{j}$ is defined in (2.4). Then for any $t_{0},t_{1},\dots,t_{\ell}\in\mathbb{R}$ and $\zeta\in\mathbb{C}$ we have
$\deg\mathrm{e}^{2\pi i(t_{0}c+t_{1}x^{1}+\dots+t_{\ell}x^{\ell}+\zeta)}=-t_{0}+t_{1}\theta_{1}+\dots+t_{\ell}\theta_{\ell}.$ (2.8)
In other words, for any given linear function $\psi(x^{1},\dots,x^{\ell};c)=t_{0}c+t_{1}x^{1}+\dots+t_{\ell}x^{\ell}$ and constant $m_{0}\in\mathbb{C}^{\times}$ we have
$\deg\left(m_{0}\mathrm{e}^{2\pi i\psi(x^{1},\dots,x^{\ell};c)}\right)=\psi(\theta_{1},\dots,\theta_{\ell};-1).$ (2.9)
The action of $W_{a}(R)$ on the affine coordinates of $V$ that is defined by (2.6) leads to a right action of $W_{a}(R)$ on $\mathscr{F}$, which is given by
$\sigma:~{}\mathscr{F}\to\mathscr{F},~{}f(x^{1},\dots,x^{\ell};c)\mapsto(f\circ\sigma)(x^{1},\dots,x^{\ell};c)$ (2.10)
for any given $\sigma\in W_{a}(R)$.
@@
The affine Weyl group $W_{a}(R)$ acts homogeneously on $\mathscr{F}$, i.e., it preserves the degrees of monomials.
@@
Let $m(x^{1},\dots,x^{\ell};c)$ be a nonzero monimial, which can be represented in the form
$m(x^{1},\dots,x^{\ell};c)=m_{0}\mathrm{e}^{2\pi i\psi(x^{1},\dots,x^{\ell};c)}$
with a linear function $\psi(x^{1},\dots,x^{\ell};c)=t_{0}c+t_{1}x^{1}+\dots+t_{\ell}x^{\ell}$ and a nonzero constant $m_{0}$. For any given translation
$\sigma:\mathbf{x}\mapsto\mathbf{x}+\sum_{r=1}^{\ell}n_{r}\alpha_{r}^{\vee}$
of $W_{a}(R)$, its action on the affine coordinates $(x^{1},\dots,x^{\ell};c)$ is given by
$(x^{1},\dots,x^{\ell};c)\mapsto(x^{1}+n_{1},\dots,x^{\ell}+n_{\ell};c),$
so the associated action on $m_{0}\mathrm{e}^{2\pi i\psi(x^{1},\dots,x^{\ell};c)}$ yields $m_{0}^{\prime}\mathrm{e}^{2\pi i\psi(x^{1},\dots,x^{\ell};c)}$, which preserves the linear function $\psi(x^{1},\dots,x^{\ell};c)$ and only changes $m_{0}$ to another constant
$m_{0}^{\prime}=m_{0}\mathrm{e}^{2\pi i(t_{1}n_{1}+\dots+t_{\ell}n_{\ell})}.$
In particular, the action by $\sigma$ preserves the degree of $m(x^{1},\dots,x^{\ell};c)$.
On the other hand, the action of any given $\sigma\in W(R)$ on the affine coordinates $(x^{1},\dots,x^{\ell};c)$ is linear, hence
$(\sigma m)(x^{1},\dots,x^{\ell};c)=m_{0}\mathrm{e}^{2\pi i(\psi\circ\sigma)(x^{1},\dots,x^{\ell};c)}.$
From the relations
$\sigma(\mathbf{0})=\mathbf{0},\quad(-1)\omega+\theta_{1}\alpha_{1}^{\vee}+\dots+\theta_{\ell}\alpha_{\ell}^{\vee}=\mathbf{0},$
it follows that $\sigma$ preserves $(\theta_{1},\dots,\theta_{\ell};-1)$. Therefore by using (2.9) we arrive at
$\deg(\sigma m)=(\psi\circ\sigma)(\theta_{1},\dots,\theta_{\ell};-1)=\psi(\theta_{1},\dots,\theta_{\ell};-1)=\deg m.$
The lemma is proved. ∎
Denote by $\mathscr{F}^{W}$ the invariant subring of $\mathscr{F}$ w.r.t. the action of $W_{a}(R)$. Let
$Y_{j}(x^{1},\dots,x^{\ell};c)=\frac{1}{N_{j}}\sum_{\sigma\in W}\mathrm{e}^{2\pi i(\omega_{j},\sigma(\mathbf{x}))},\quad j=1,\dots,\ell,$ (2.11)
where $N_{j}=\#\{\sigma\in W\mid\sigma(\omega_{j})=\omega_{j}\}$ and $\mathbf{x}=c\omega+x^{1}\alpha_{1}^{\vee}+\dots+x^{\ell}\alpha_{\ell}^{\vee}$.
@@
The functions $Y_{j}$ belong to $\mathscr{F}^{W}$ and are homogeneous of degree 00.
@@
Let us proceed to consider generators of the $W_{a}(R)$-invariant ring $\mathscr{F}^{W}$. Since $\mathscr{F}^{W}$ is too large to consider its generators, we are to introduce a smaller subring of $\mathscr{F}$ which is named the $\lambda$-Fourier polynomial ring.
Denote
@@
In particular, if we take $\omega=\omega_{k}$, then $\kappa=(\omega_{k},\alpha_{k})$ and $\lambda=\mathrm{e}^{-2\pi ic(\omega_{k},\alpha_{k})}$. Define the $\lambda$-Fourier polynomial ring $\mathscr{A}$ as a subring of $\mathscr{F}$ by
@@
and their linear combinations. We denote by $\mathscr{A}^{W}=\mathscr{A}\cap\mathscr{F}^{W}$ the $W_{a}(R)$-invariant $\lambda$-Fourier polynomial ring. Then we have $\lambda=\mathrm{e}^{-2\pi i\kappa c}\in\mathscr{A}^{W}$.
@@
$y^{j}(\mathbf{x})=\mathrm{e}^{-2\pi i\theta_{j}c}~{}Y_{j}(\mathbf{x}),~{}j=1,\dots,\ell.$ (2.15)
From Proposition 2.2 it follows that $y^{j}$ is homogeneous of degree $\theta_{j}$.
@@
$-2\pi i\theta_{j}c+2\pi i(\omega_{j},\sigma(\mathbf{x}))~{}=$ $~{}-2\pi ic(\omega_{j},\omega)+2\pi i(\sigma^{*}(\omega_{j}),\mathbf{x})$
$=$ $~{}-2\pi ic(\omega_{j}-\sigma^{*}(\omega_{j}),\omega)+2\pi i\sum_{r=1}^{\ell}x^{r}(\sigma^{*}(\omega_{j}),\alpha_{r}^{\vee}).$ (2.17)
Since $\sigma^{*}\in W(R)$, there exist non-negative integers $n_{1},\dots,n_{\ell}$ such that [6]
$\sigma^{*}(\omega_{j})=\omega_{j}-\sum_{s=1}^{\ell}n_{s}\alpha_{s},$ (2.18)
@@
$(\omega_{j}-\sigma^{*}(\omega_{j}),\omega)=\sum_{s=1}^{\ell}n_{s}\left(\alpha_{s},\omega\right)=m\kappa,$ (2.19)
@@
$\left(\sigma^{*}(\omega_{j}),\alpha_{r}^{\vee}\right)=(\omega_{j},\alpha_{r}^{\vee})-\sum_{s=1}^{\ell}n_{s}(\alpha_{s},\alpha_{r}^{\vee})=\delta_{jr}-\sum_{s=1}^{\ell}n_{s}c_{sr},$ (2.20)
@@
We consider the case where the root system $R=A_{2}$ and the weight $\omega=\omega_{2}$. We take the simple roots
$\alpha_{1}=e_{1}-e_{2},\quad\alpha_{2}=e_{2}-e_{3}$
of the root system $A_{2}$ in the hyperplane $V$ of the Euclidean space $\mathbb{R}^{3}$ spanned by $\alpha_{1},\alpha_{2}$. We have
@@
The affine coordinates $(x^{1},x^{2};c)$ on $V$ are given by
@@
The Weyl group $W$ acting on $V$ is generated by the reflections
$\sigma_{j}(\mathbf{x})=\mathbf{x}-(\alpha^{\vee}_{j},\mathbf{x})\alpha_{j},\quad j=1,2,\ \mathbf{x}\in V.$
@@
Here $\kappa=1$, hence
$\lambda=\mathrm{e}^{-2\pi ic}.$
Then we have
@@
In the next subsection we are to show that $y^{1},\dots,y^{\ell}$ and $\lambda$ generate $\mathscr{A}^{W}$.
2.3. The $\varphi_{\lambda}$-transformation and the null transformation
The parameter $\lambda$ of the $\lambda$-Fourier polynomials $y^{1},\dots,y^{\ell}$ is defined by (2.13), so it lies on the unit circle. However, since $y^{j}$ depend polynomially on $\lambda$, the definition of $y^{j}$ can be extended naturally to any $\lambda\in\mathbb{C}$, including $\lambda=0$.

### v3

@@ Abstract.
We present an approach to construct a class of generalized Frobenius manifold structures on the orbit spaces of affine Weyl groups, and prove that their monodromy groups are parabolic subgroups of the associated affine Weyl groups.
@@
Generalized Frobenius manifold, Affine Weyl group, Flat pencil of metrics, Monodromy Group.
@@ 1. Introduction
The notion of Frobenius manifold was introduced by Dubrovin in the early 1990s in his study on the relation of integrable systems and 2-dimensional topological field theories [8, 9]. This geometric structure and its generalizations play a central role in topological field theory, Gromov-Witten theory, singularity theory and in the theory of integrable systems (see [1, 2, 3, 4, 5, 9, 13, 15, 18, 19, 22, 24] and the references therein).
@@
In this paper, we construct on the orbit spaces of affine Weyl groups a class of generalized Frobenius manifold structures. These generalized structures satisfy all the axioms of Dubrovin’s definition of Frobenius manifolds except for the flatness of the unit vector fields. The simplest example in this class corresponds to the affine Weyl group of type $A_{1}$, which naturally appears in the study of Hodge integrals and their remarkable connections to two important discrete integrable systems: the Volterra hierarchy and the $q$-deformed KdV hierarchy — both fundamental in soliton theory[21, 22].
@@
Our approach parallels the construction of a Frobenius manifold structure on the orbit space of a Coxeter group or on a certain extension of an affine Weyl group [9, 12, 14, 25, 26, 28, 29]. For these cases, a flat pencil of metrics is obtained by projecting the Euclidean metric (on the space where the group acts) onto the orbit space, based on which the Frobenius manifold structure is constructed. The key to these constructions lies in the linear dependence of both the contravariant components of the metric and its Levi-Civita connection on a distinguished coordinate of the orbit space — this special coordinate naturally gives rise to the flat unit vector field of the Frobenius manifold. However, for the generalized Frobenius manifolds that we are considering, the unit vector field is not flat, implying the absence of such a special coordinate. To overcome this, we proceed as follows: Given an affine Weyl group $W_{a}$ acting on an $\ell$-dimensional Euclidean space $V$, we introduce a parameter $\lambda$ into the invariant Fourier polynomials, yielding the coordinates $z^{1},\dots,z^{\ell}$ on the orbit space. The projected contravariant metric $\left(g_{\lambda}^{\alpha\beta}(z)\right)$ then depends on $\lambda$. Our approach is to choose the invariant $\lambda$-Fourier polynomials such that $\left(g_{\lambda}^{\alpha\beta}(z)\right)$ and the Christoffel symbols of their Levi-Civita connections depend linearly on $\lambda$. We prove in this paper that as long as this linear dependence holds, there exists a corresponding generalized Frobenius manifold structure on a dense open subset of the orbit space.
@@
In the present paper, we detail this construction, study properties of the resulting generalized Frobenius manifolds including their monodromy groups, and provide some explicit examples. Further examples will appear in our subsequent paper [20]. The paper is organized as follows. In Sect. 2, we introduce the invariant $\lambda$-Fourier polynomial ring and study its properties. In Sect. 3, we construct generalized Frobenius manifold structures on orbit spaces of affine Weyl groups, and summarize the main results in Theorem 3.13. In Sect. 4, we analyze properties of the monodromy groups of these generalized Frobenius manifolds. In Sect. 5, we present some representative examples. In Sect. 6, we conclude the paper with remarks and further directions.
@@ 2.1. Affine coordinates
Let $R$ be an irreducible reduced root system in an $\ell$-dimensional Euclidean space $V$ with inner product (,)($\cdot\,$$\cdot$), and $\alpha_{1},\dots,\alpha_{\ell}$ be a fixed basis of simple roots. Denote by $\alpha_{1}^{\vee},\dots,\alpha_{\ell}^{\vee}$ the corresponding coroots given by
@@
where $c\in\mathbb{R}$ is a fixed parameter. We will also use the following expressions of vectors in $V$:
@@
$\phi\colon\mathbf{x}\mapsto\phi_{0}(\mathbf{x})+\sum_{j=1}^{\ell}k_{j}\alpha_{j}^{\vee},\quad\phi_{0}\in W(R),\ k_{1},\dots,k_{\ell}\in\mathbb{Z},$ (2.5)
@@
2.2. The invariant $\lambda$-Fourier polynomial ring and basic generators
In this subsection, we are to introduce the $\lambda$-Fourier polynomial ring $\mathscr{A}$ and its invariant subring $\mathscr{A}^{W}$ w.r.t. the action of the affine Weyl group $W_{a}(R)$, and define the set of basic generators $\{y^{1},\dots,y^{\ell}\}$. To this end, we first introduce a bigger ring $\mathscr{F}$, called the Fourier function ring, as follows:
$\mathscr{F}=\mathrm{span}_{\mathbb{C}}\left\{f(x^{1},\dots,x^{\ell};c)=\mathrm{e}^{2\pi i(t_{0}c+t_{1}x^{1}+\dots+t_{\ell}x^{\ell})}\mid t_{0},t_{1},\dots,t_{\ell}\in\mathbb{C}\right\}.$
For any $f\in\mathscr{F}$, we can uniquely write it in the form $f(\mathbf{x};c)$. Then we have a $W_{a}(R)$-action on $\mathscr{F}$ given by
$(\phi^{*}f)(\mathbf{x};c)=f(\phi(\mathbf{x});c),\quad\phi\in W_{a}(R).$ (2.6)
Note that if we write $f$ as a $f(\hat{\mathbf{x}};c)$, then the action is given by
$(\phi^{*}f)(\hat{\mathbf{x}};c)=(\phi^{*}f)(\mathbf{x}-c\omega;c)=f(\phi(\mathbf{x})-c\omega;c),\quad\phi\in W_{a}(R),$ (2.7)
which in general does not coincide with $f(\phi(\hat{\mathbf{x}});c)$. We introduce a $\mathbb{C}$-gradation on $\mathscr{F}$ by the assignment
@@ 2.2. The invariant $\lambda$-Fourier polynomial ring and basic generators
where $\theta_{j}$ is defined in (2.4). Then for any given linear function
$\psi(\mathbf{x};c)=t_{1}(x^{1}+\theta_{1}c)+\dots+t_{\ell}(x^{\ell}+\theta_{\ell}c)+t_{0}c,\quad t_{1},\dots,t_{\ell}\in\mathbb{C}$ (2.9)
and constant $A_{0}\in\mathbb{C}^{\times}$, we have
$\deg\left(A_{0}~\mathrm{e}^{2\pi i\psi(\mathbf{x};c)}\right)=\psi(\mathbf{0};-1).$ (2.10)
Here $\mathbf{0}$ is the zero vector in $V$.
@@
The affine Weyl group $W_{a}(R)$ acts quasi-homogeneously on $\mathscr{F}$, i.e., it preserves the degrees of monomials.
@@
Let $f(\mathbf{x};c)\in\mathscr{F}$ be a nonzero monomial, which can be represented in the form
$f(\mathbf{x};c)=A_{0}\mathrm{e}^{2\pi i\psi(\mathbf{x};c)}$
with a linear function $\psi$ of the form (2.9) and a nonzero constant $A_{0}$. On one hand, for any given translation
$\phi\colon\mathbf{x}\mapsto\mathbf{x}+\alpha,\quad\alpha=\sum_{r=1}^{\ell}n_{r}\alpha_{r}^{\vee}$
of $W_{a}(R)$, its action on $f$ is given by
$(\phi^{*}f)(\mathbf{x};c)=A_{0}\mathrm{e}^{2\pi i\psi(\mathbf{x}+\alpha;c)}=\mathrm{e}^{2\pi i\psi(\alpha;0)}f(\mathbf{x};c).$
Therefore, $\phi^{*}f$ and $f$ merely differ by a nonzero factor, hence the action by $\phi$ preserves the degree.
On the other hand, for any $\phi\in W(R)$ is given by
$\deg(\phi^{*}f)=\deg\left(A_{0}~\mathrm{e}^{2\pi i\psi(\phi(\mathbf{x});c)}\right)=\psi(\phi(\mathbf{0});c).$
Since $\phi$ is linear, we have $\phi(\mathbf{0})=\mathbf{0}$, hence $\deg(\phi^{*}f)=\deg f$. The lemma is proved. ∎
Denote by $\mathscr{F}^{W}$ the invariant subring of $\mathscr{F}$ w.r.t. the action of $W_{a}(R)$. We introduce the invariant functions $Y_{1},\dots,Y_{\ell}\in\mathscr{F}^{W}$ by
$Y_{j}(x^{1},\dots,x^{\ell};c)=\frac{1}{N_{j}}\sum_{\sigma\in W(R)}\mathrm{e}^{2\pi i(\omega_{j},\sigma(\mathbf{x}))},\quad j=1,\dots,\ell,$ (2.11)
where $N_{j}=\#\{\sigma\in W(R)\colon\sigma(\omega_{j})=\omega_{j}\}$ and $\mathbf{x}=c\omega+x^{1}\alpha_{1}^{\vee}+\dots+x^{\ell}\alpha_{\ell}^{\vee}$.
@@
The functions $Y_{j}$ belong to $\mathscr{F}^{W}$ and are quasi-homogeneous of degree 0.
@@
Let us proceed to introduce the $\lambda$-Fourier polynomial ring $\mathscr{A}$ and its invariant subring $\mathscr{A}^{W}$ w.r.t. the action of $W_{a}(R)$. Denote
@@
Define the $\lambda$-Fourier polynomial ring $\mathscr{A}$ as a subring of $\mathscr{F}$ by
@@
and their linear combinations. We call the subring of $\mathscr{A}$ defined by
$\mathscr{A}^{W}=\mathscr{A}\cap\mathscr{F}^{W}$
the invariant $\lambda$-Fourier polynomial ring. Then we have $\lambda=\mathrm{e}^{-2\pi i\kappa c}\in\mathscr{A}^{W}$. Note that $W_{a}(R)$ does not act on $\mathscr{A}$, since results of the action may involve negative power terms of $\lambda$, which are not in $\mathscr{A}$.
@@
$y^{j}(\mathbf{x};c)=\mathrm{e}^{-2\pi i\theta_{j}c}~Y_{j}(\mathbf{x}),~j=1,\dots,\ell.$ (2.15)
It follows from Proposition 2.2 that $y^{j}$ is quasi-homogeneous of degree $\theta_{j}$.
@@
$-2\pi i\theta_{j}c+2\pi i(\omega_{j},\sigma(\mathbf{x}))~=$ $~-2\pi ic(\omega_{j},\omega)+2\pi i(\sigma^{-1}(\omega_{j}),\mathbf{x})$
$=$ $~-2\pi ic(\omega_{j}-\sigma^{-1}(\omega_{j}),\omega)+2\pi i\sum_{r=1}^{\ell}x^{r}(\sigma^{-1}(\omega_{j}),\alpha_{r}^{\vee}).$ (2.17)
Since $\sigma^{-1}\in W(R)$, there exist non-negative integers $n_{1},\dots,n_{\ell}$ such that [6]
$\sigma^{-1}(\omega_{j})=\omega_{j}-\sum_{s=1}^{\ell}n_{s}\alpha_{s},$ (2.18)
@@
$(\omega_{j}-\sigma^{-1}(\omega_{j}),\omega)=\sum_{s=1}^{\ell}n_{s}\left(\alpha_{s},\omega\right)=m\kappa,$ (2.19)
@@
$\left(\sigma^{-1}(\omega_{j}),\alpha_{r}^{\vee}\right)=(\omega_{j},\alpha_{r}^{\vee})-\sum_{s=1}^{\ell}n_{s}(\alpha_{s},\alpha_{r}^{\vee})=\delta_{jr}-\sum_{s=1}^{\ell}n_{s}c_{sr},$ (2.20)
@@
Theorem 2.4.
We have $\mathscr{A}^{W}=\mathbb{C}[y^{1},\dots,y^{\ell};\lambda]$.
A proof of this theorem will be given in the next subsection. We call $\{y^{1},\dots,y^{\ell}\}$ the set of basic generators of $\mathscr{A}^{W}$.
@@
Let $e_{1},e_{2},e_{3}$ be an the orthonormal basis of the Euclidean space $\mathbb{R}^{3}$, and $V$ be the hyperplane spanned by
$\alpha_{1}=e_{1}-e_{2},\quad\alpha_{2}=e_{2}-e_{3}.$
Then we can take $\alpha_{1},\alpha_{2}$ as a basis of simple roots for the root system $R$ of type $A_{2}$ in $V$. We have
@@
Let us take $\omega=\omega_{2}$, then the affine coordinates $(x^{1},x^{2};c)$ on $V$ are given by
@@
Since $\kappa=1$, we have $\lambda=\mathrm{e}^{-2\pi ic}$, and the basic generators are given by
@@
2.3. The $\varphi_{\lambda}$-transformation and the null-transformation
The parameter $\lambda$ of the $\lambda$-Fourier polynomials $y^{1},\dots,y^{\ell}$ is defined by (2.13), so it lies on the unit circle. However, since $y^{j}$ depend polynomially on $\lambda$, the definition of $y^{j}$ can be extended naturally to any $\lambda\in\mathbb{C}$, including $\lambda=0$ in particular.